# An Appropriate Probability Model for the Bell Experiment

### v2

@@ 6.2 Non-separability
$\sum_{x,y}{x.y.p(x,y|\alpha,\beta)}~=~\sum_{\lambda\in\Lambda}\Big(\sum_{x}{x.p_{1}(x|\alpha,\lambda)}\cdot\sum_{y}y.p_{2}(y|\beta,\lambda)\Big)\cdot\rho(\lambda)~,$ (6.5)
@@
We use this, by choosing the elements $a_{0},a_{1},b_{0},b_{1}$ equal to Tsirelson’s angles (4.6) and we define
@@
Note that $Q_{\lambda},R_{\lambda},T_{\lambda},$ and $S_{\lambda},$ are all in [1,1][-1,1].
So, by the assumption of separability, we obtain:
$\begin{array}[]{lll}\mathbb{E}[XY|A=a_{0},B=b_{0}]&=&\sum_{\lambda}Q_{\lambda}S_{\lambda}\rho(\lambda),~\mathbb{E}[XY|A=a_{0},B=b_{1}]=\sum_{\lambda}Q_{\lambda}T_{\lambda}\rho(\lambda)\\[2.84526pt] \mathbb{E}[XY|A=a_{1},B=b_{0}]&=&\sum_{\lambda}R_{\lambda}S_{\lambda}\rho(\lambda),~\mathbb{E}[XY|A=a_{1},B=b_{1}]=\sum_{\lambda}R_{\lambda}T_{\lambda}\rho(\lambda)~.\end{array}$ (6.8)
@@
Equation (6.10) is in clear contradiction with (6.13). This contradiction proves that the probability measure of our probability model is not Bell separable. Our proof here for non-separability is a formalization of sketch offered by [8]. A different proof without the CHSH-inequality is given in Appendix A, where we also prove that if separability would hold, the process is deterministic, i.e. the probability measure $\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta]\in\{0,1\}$.
In summary we conclude that there does not exist a hidden variable that makes our probability separable. Note that we did not assume any structure of the hidden variable.
7 Conclusion
@@
This paper proposes an appropriate probability model for the Bell experiment. The outcome space is defined as the cartesian product space of random variables $A$ and $B$ (detector settings) and $X$ and $Y$, the corresponding measurement outcomes. It has only two simultaneously observable detector settings per measurement and therefore does not need the assumption of realism. The probability measure $p(x,y,a,b)$ is derived from quantum mechanics and described by Tables 2 and 3.
From this probability measure it follows that the conditional expectation $\mathbb{E}[X\!\cdot\!Y\;\big|\;A\!=\!\alpha,B\!=\!\beta]$ corresponds to Born expectation $\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}$. The Bell contradiction arises from an invalid summation of conditional expectations with different conditions (i.e. different detector settings). That is, the step from (4.3) to (4.4) is not allowed for conditional expectations.
Aspect et al [4] presented their experimental results as “the greatest violation of generalised Bell’s inequalities ever achieved”. However, our probability model of the (generalized) EPR/Bell/CHSH experiment implies that the notion of “violation of Bell’s inequality” has no meaning in quantum mechanics. Their measurements do confirm quantum mechanical expectation (3.9) with high precision.
@@
In Section 6.2 we showed that our probability model cannot be extended with a hidden variable such that its probability measure is separable in the Bell sense. [^7] The proof in Section 6.2 makes use of the CHSH inequality. The alternative proof in Appendix 7 does not. This non-separability of the probability measure supports Bell’s [19] conclusion “If [a hidden-variable theory] is local it will not agree with quantum mechanics, and if it agrees with quantum mechanics it will not be local.”
Acknowledgements
@@
The authors thank Richard Gill (Leiden University), Andrea Fiore, Servaas Kokkelmans, and Henk Nijmeijer (all Technical University Eindhoven), and Sven Jandura, (University of Strasbourg) for providing helpful feedback on an earlier version of this paper.
@@ Appendix A Alternative proof for non-separability
Table (2) implies that the outcomes of single-sided experiments must satisfy
@@
The extended outcome space is depicted in Table 4, together with the assumed factorizations in terms of $p$ and $q$ for the measurement outcomes, given detector settings $a$ and $b$. Note that the right column contains probability densities in stead of probabilities.
@@
1 1 $\frac{1}{2}\sin^{2}(a\!-\!b)$ $p(a,\lambda).(1-q(b,\lambda))$
-1 1 $\frac{1}{2}\cos^{2}(a\!-\!b)$ $(1-p(a,\lambda)).(1-q(b,\lambda))$
1 -1 $\frac{1}{2}\cos^{2}(a\!-\!b)$ $p(a,\lambda).q(b,\lambda)$
-1 -1 $\frac{1}{2}\sin^{2}(a\!-\!b)$ $(1-p(a,\lambda)).q(b,\lambda)$
@@
For specific outcome (+1,1)(+1,-1) Eqn (3.8) states that $\mathbb{P}[+1,-1\;\big|\;a,b]=\frac{1}{2}\cos^{2}(a\!-\!b)$, hence Bell-separability implies:
$\int_{\Lambda}p(a,\lambda)\cdot q(b,\lambda)\cdot{\rm d}\rho(\lambda)=\frac{1}{2}\cos^{2}(a\!-\!b)~.$ (A.3)
@@
Step 1: rewrite the above conditions in terms of $\tilde{p},\tilde{q}$, using $\tilde{p}=2p\!-\!1$, $\tilde{q}=2q\!-\!1$, and $\cos(2(a\!-\!b))=2\cos^{2}(a\!-\!b)\!-\!1$:
$\begin{array}[]{lll}\tilde{p},\tilde{q}:[0,\pi]\times\Lambda\to[-1,1]~,&&(i)\\[4.2679pt] \int_{\Lambda}\tilde{p}(a,\lambda){\rm d}\rho(\lambda)=0,~~~~\int_{\Lambda}\tilde{q}(a,\lambda){\rm d}\rho(\lambda)=0~,&&(ii)\\[4.2679pt] \int_{\Lambda}\tilde{p}(a,\lambda)\tilde{q}(b,\lambda){\rm d}\rho(\lambda)=\cos\big(2(a\!-\!b)\big)~.&&(iii)\end{array}$ (A.4)
Step 2: demonstrate that $\tilde{p}=\tilde{q}~$ and that $\tilde{p}(a,\lambda)\in\{-1,1\}$. These properties follow from taking a=ba\!=\!b and inspecting of the following integrals:
$\int_{\Lambda}(1+\tilde{p}(a,\lambda))(1-\tilde{q}(a,\lambda)){\rm d}\rho(\lambda)=0,~~~~\int_{\Lambda}(1+\tilde{q}(a,\lambda))(1-\tilde{p}(a,\lambda)){\rm d}\rho(\lambda)=0~.$ (A.5)
@@
$1+\tilde{p}(a,\lambda)-\tilde{q}(a,\lambda)-\tilde{p}(a,\lambda)\tilde{q}(a,\lambda)=0,~~~~1+\tilde{q}(a,\lambda)-\tilde{p}(a,\lambda)-\tilde{p}(a,\lambda)\tilde{q}(a,\lambda)=0~,$ (A.6)
which leads to $\tilde{p}(a,\lambda)=\tilde{q}(a,\lambda)~.$ As a result of this equality $\int_{\Lambda}(1-\tilde{p}^{2}(a,\lambda)){\rm d}\rho(\lambda)=0~$, and $\tilde{p}(a,\lambda)$ can necessarily only assume values in $\{-1,1\}$. [^8] Assuming that such $\tilde{p}$ and $\tilde{q}$ exist, this constraint implies deterministic measurement outcomes.
@@
$\big\{\,a\mapsto\frac{e^{2{\rm i}ka}}{\sqrt{\pi}}\big\}_{k\in\mathbb{Z}}~~~~~~~~~~~~\big\{(a,b)\mapsto\frac{e^{2{\rm i}ka}e^{2{\rm i}\ell b}}{\pi}\big\}_{(k,\ell)\in\mathbb{Z}^{2}}~.$ (A.7)
@@
$\tilde{p}(a,\lambda)=\sum_{k\in\mathbb{Z}}\,c_{k}(\lambda)\frac{e^{2{\rm i}ka}}{\sqrt{\pi}},~~~\text{where }~~~c_{k}(\lambda)=\int_{0}^{\pi}\tilde{p}(a,\lambda)\frac{e^{-2{\rm i}ka}}{\sqrt{\pi}}{\rm d}a~,$ (A.8)
@@
$\cos\big(2(a\!-\!b)\big)=\frac{\pi}{2}\frac{e^{2{\rm i}a}e^{-2{\rm i}b}}{\pi}+\frac{\pi}{2}\frac{e^{-2{\rm i}a}e^{2{\rm i}b}}{\pi}$ (A.9)
@@
$\begin{array}[]{lll}\int_{\Lambda}\tilde{p}(a,\lambda){\rm d}\rho(\lambda)=\sum_{k\in\mathbb{Z}}\Big(\int_{\Lambda}c_{k}(\lambda){\rm d}\rho(\lambda)\Big)\frac{e^{2{\rm i}ka}}{\sqrt{\pi}}=0,&&(ii)\\[5.69054pt] \sum_{(k,\ell)\in\mathbb{Z}^{2}}\Big(\int_{\Lambda}c_{k}(\lambda)c_{l}(\lambda){\rm d}\rho(\lambda)\Big)\frac{e^{2{\rm i}ka}e^{2{\rm i}\ell b}}{\pi}~=~\frac{\pi}{2}\frac{e^{2{\rm i}a}e^{-2{\rm i}b}}{\pi}+\frac{\pi}{2}\frac{e^{-2{\rm i}a}e^{2{\rm i}b}}{\pi}.&&(iii)\end{array}$ (A.10)
@@
$\begin{array}[]{lll}c_{0}(\lambda)\in\mathbb{R},~~~\forall k\in\mathbb{Z}:&\int_{\Lambda}c_{k}(\lambda){\rm d}\rho(\lambda)&=0.\\[5.69054pt] &\int_{\Lambda}c_{1}(\lambda)c_{-1}(\lambda){\rm d}\rho(\lambda)&=\frac{\pi}{2}\\[5.69054pt] \forall(k,\ell)\in\mathbb{Z}^{2}\backslash\{(1,-1),(-1,1)\}:&\int_{\Lambda}c_{k}(\lambda)c_{l}(\lambda){\rm d}\rho(\lambda)&=0\\[5.69054pt] \forall k\in\mathbb{Z}\backslash\{-1,1\}:&\int_{\Lambda}c_{k}(\lambda)c_{-k}(\lambda){\rm d}\rho(\lambda)&=0~.\end{array}$ (A.11)
@@
$\tilde{p}(a,\lambda)=\dfrac{1}{\sqrt{\pi}}\big(c_{1}(\lambda)e^{2{\rm i}a}+c_{-1}(\lambda)e^{-2{\rm i}a}\big)~=~\dfrac{2|c_{1}(\lambda)|}{\sqrt{\pi}}\cos\big(2a+\alpha(\lambda)\big)~,$ (A.13)
where $\alpha(\lambda)$ is given by $c_{1}=|c_{1}(\lambda)|e^{-i\alpha(\lambda)}$. The values of $c_{\pm 1}(\lambda)$ depend on the choice for $\Lambda$. For example, if $\Lambda=[0,1]\subset\mathbb{R}$ then condition (A.12) is met by $c_{\pm 1}(\lambda)=\sqrt{\frac{\pi}{2}}e^{\pm 2\pi{\rm i}\lambda}$. Note that $\tilde{p}(a,\lambda)$ assumes many more values than just $\pm 1$ as was established in Step 2. With this contradiction in the derived properties of $\tilde{p}(a,\lambda)$ we conclude that $p(x,y,a,b)$ is not Bell separable.

### v3

@@ 6.2 Bell Separability
$\sum_{x,y}{x\cdot~y\cdot~p(x,y|\alpha,\beta)}~=~\sum_{\lambda\in\Lambda}\Big(\sum_{x}{x\cdot~p_{1}(x|\alpha,\lambda)}\cdot\sum_{y}y\cdot~p_{2}(y|\beta,\lambda)\Big)\cdot\rho(\lambda)~,$ (6.9)
@@
Next, by choosing the elements $a_{0},a_{1},b_{0},b_{1}$ equal to Tsirelson’s angles (4.6) and we define
@@
Note that $Q_{\lambda},R_{\lambda},T_{\lambda},$ and $S_{\lambda},$ are all in [1,1][-1,1]. Their substitution in (6.10) results in.
$\begin{array}[]{lll}\mathbb{E}[XY|A\!=\!a_{0},B\!=\!b_{0}]&=&\sum_{\lambda}Q_{\lambda}S_{\lambda}~\rho(\lambda)\\[2.84526pt] \mathbb{E}[XY|A\!=\!a_{0},B\!=\!b_{1}]&=&\sum_{\lambda}Q_{\lambda}T_{\lambda}~\rho(\lambda)\\[2.84526pt] \mathbb{E}[XY|A\!=\!a_{1},B\!=\!b_{1}]&=&\sum_{\lambda}R_{\lambda}S_{\lambda}~\rho(\lambda)\\[2.84526pt] \mathbb{E}[XY|A\!=\!a_{1},B\!=\!b_{1}]&=&\sum_{\lambda}R_{\lambda}T_{\lambda}~\rho(\lambda)~.\end{array}$ (6.12)
@@
Equation (6.14) is in clear contradiction with (6.17). This contradiction proves that the probability measure is not Bell separable. This proof is a formalization of a sketch offered by [8]. An alternative proof without the use of the CHSH-inequality is given in Appendix A. In summary we conclude that there does not exist a hidden variable that makes our probability measure separable. Note that we did not assume any structure of the hidden variable, except measurement independence (6.3).
6.3 Determinism
The model is called deterministic if for all $\lambda\in\Lambda$ [9]
$\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!a,B\!=\!b,L\!=\!\lambda]~\in~\{0,1\}~.$ (6.18)
For a model that is (statistically) local (6.4), Bell separability (6.5) is equivalent to determinism:
$\cal S\land\cal L~\Leftrightarrow~\cal D\land\cal$ L~, (6.19)
where $\cal S,\cal$ L, and $\cal D$ denote the properties separable, statistically local, and deterministic, respectively. The immediate implication is
$\neg\cal S~\Rightarrow~\neg\cal D\lor\neg\cal$ L~. (6.20)
In words, the proposed probability model is not separable, and therefore it is non-deterministic or non-local (or both).
The proof of (6.19) is presented in two parts. First we focus on $\cal S\land\cal L~\Rightarrow~\cal D$, which makes use of the following lemma.
$\begin{array}[]{llll}&p_{1}(x|\alpha,\beta,\lambda)\in\{0,1\}~\land~p_{2}(y|\alpha,\beta,\lambda)\in\{0,1\}\\[2.84526pt] \Rightarrow&p(x,y|\alpha,\beta,\lambda)~=~p_{1}(x|\alpha,\beta,\lambda)\cdot p_{2}(y|\alpha,\beta,\lambda)~\in\{0,1\}~,\end{array}$ (6.21)
where $p$, $p_{1}$, and $p_{2}$ are obvious generalizations of those in (6.5). From $p_{1}()$ we conclude that for only one $x$ we have $p_{1}(x|\alpha,\beta,\lambda)\neq 0$ and similarly for $p_{2}()$. Hence, for only one pair $x,y$ it holds that $p(x,y|\alpha,\beta,\lambda)\neq 0$. For that pair the probability must be 1.
Given this lemma, we prove $\cal S\land\cal L~\Rightarrow~\cal D$~ by
[…]
The second part of the proof makes use of a similar lemma.
$\begin{array}[]{lll}&p(x,y|\alpha,\beta,\lambda)~\in~\{0,1\}\\[2.84526pt] \Rightarrow&p_{1}(x|\alpha,\beta,\lambda)\in\{0,1\}~\land~p_{2}(y|\alpha,\beta,\lambda)\in\{0,1\}\\[2.84526pt] &~\land~p(x,y|\alpha,\beta,\lambda)~=~p_{1}(x|\alpha,\beta,\lambda)\cdot p_{2}(y|\alpha,\beta,\lambda)~,\end{array}$ (6.23)
because there can be only one $x,y$ for which $p(x,y|\alpha,\beta,\lambda)=1$.
Given this lemma, we prove $\cal D\land\cal L~\Rightarrow~\cal S$~ by
[…]
That is, given $\cal D\land\cal$ L, the probability model is Bell-separable (6.5).
7 Conclusions
@@
This paper proposes a probability model for the Bell experiment. The outcome space is defined as the cartesian product space of random variables $A$ and $B$ (detector settings) and $X$ and $Y$, the corresponding measurement outcomes. It has only two simultaneously observable detector settings per measurement and therefore does not need the assumption of realism. The probability measure $p(x,y,a,b)$ is derived from quantum mechanics and described by Tables 2 and 3.
[…]
@@
In Section 6 we extend the probability model with a hidden variable. This extended probability model is proved to be non-separable in the Bell sense. (The proof in Section 6.2 makes use of the CHSH inequality. The alternative proof in Appendix A does not.) This non-separability of the probability measure supports Bell’s [16] conclusion “If [a hidden-variable theory] is local it will not agree with quantum mechanics, and if it agrees with quantum mechanics it will not be local”. Furthermore, non-separability implies that the model is non-deterministic or non-local (or both).
@@
Acknowledgments: The authors thank Richard Gill (Leiden University), Andrea Fiore, Servaas Kokkelmans, and Henk Nijmeijer (all Technical University Eindhoven), and Sven Jandura, (University of Strasbourg) for providing helpful feedback on an earlier version of this paper. The feedback of the anonymous reviewers helped us to greatly improve the presentation of the material in Section 6.
@@ Appendix A Alternative Proof for Non-Separability
Table 2 implies that the outcomes of single-sided experiments must satisfy
@@
With this notation Table 4 is adapted into Table 5.
@@
1 1 $\frac{1}{2}\sin^{2}(a\!-\!b)$ $p(a,\lambda)\cdot(1-q(b,\lambda))$
-1 1 $\frac{1}{2}\cos^{2}(a\!-\!b)$ $(1-p(a,\lambda))\cdot(1-q(b,\lambda))$
1 -1 $\frac{1}{2}\cos^{2}(a\!-\!b)$ $p(a,\lambda)\cdot q(b,\lambda)$
-1 -1 $\frac{1}{2}\sin^{2}(a\!-\!b)$ $(1-p(a,\lambda))\cdot q(b,\lambda)$
@@
For specific outcome (+1,1)(+1,-1) Equation (3.11) states that $\mathbb{P}[+1,-1\;\big|\;\alpha,\beta]=\frac{1}{2}\cos^{2}(\alpha\!-\!\beta)$, hence Bell-separability implies:
$\int_{\Lambda}p(\alpha,\lambda)\cdot q(\beta,\lambda)\cdot{\rm d}\rho(\lambda)=\frac{1}{2}\cos^{2}(\alpha\!-\!\beta)~.$ (A.3)
@@
Step 1: rewrite the above conditions in terms of $\tilde{p},\tilde{q}$, using $\tilde{p}=2p\!-\!1$, $\tilde{q}=2q\!-\!1$, and $\cos(2(\alpha\!-\!\beta))=2\cos^{2}(\alpha\!-\!\beta)\!-\!1$:
$\begin{array}[]{lll}\tilde{p},\tilde{q}:[0,\pi]\times\Lambda\to[-1,1]~,&&(i)\\[4.2679pt] \int_{\Lambda}\tilde{p}(\alpha,\lambda){\rm d}\rho(\lambda)=0,~~~~\int_{\Lambda}\tilde{q}(\alpha,\lambda){\rm d}\rho(\lambda)=0~,&&(ii)\\[4.2679pt] \int_{\Lambda}\tilde{p}(\alpha,\lambda)\tilde{q}(\beta,\lambda){\rm d}\rho(\lambda)=\cos\big(2(\alpha\!-\!\beta)\big)~.&&(iii)\end{array}$ (A.4)
Step 2: demonstrate that $\tilde{p}=\tilde{q}~$ and that $\tilde{p}(\alpha,\lambda)\in\{-1,1\}$. These properties follow from taking a=ba\!=\!b and inspecting of the following integrals:
$\int_{\Lambda}(1+\tilde{p}(\alpha,\lambda))(1-\tilde{q}(\alpha,\lambda)){\rm d}\rho(\lambda)=0,~~~~\int_{\Lambda}(1+\tilde{q}(\alpha,\lambda))(1-\tilde{p}(\alpha,\lambda)){\rm d}\rho(\lambda)=0~.$ (A.5)
@@
$1+\tilde{p}(\alpha,\lambda)-\tilde{q}(\alpha,\lambda)-\tilde{p}(\alpha,\lambda)\tilde{q}(\alpha,\lambda)=0,~~~~(\alpha,\lambda)-\tilde{p}(\alpha,\lambda)-\tilde{p}(\alpha,\lambda)\tilde{q}(\alpha,\lambda)=0~,$ (A.6)
which leads to $\tilde{p}(\alpha,\lambda)=\tilde{q}(\alpha,\lambda)~.$ As a result of this equality $\int_{\Lambda}(1-\tilde{p}^{2}(\alpha,\lambda)){\rm d}\rho(\lambda)=0~$, and $\tilde{p}(\alpha,\lambda)$ can necessarily only assume values in $\{-1,1\}$. (Assuming that such $\tilde{p}$ and $\tilde{q}$ exist, this constraint implies deterministic measurement outcomes.)
@@
$\big\{\,\alpha\mapsto\frac{e^{2{\rm i}k\alpha}}{\sqrt{\pi}}\big\}_{k\in\mathbb{Z}}~~~~~~~~~~~~\big\{(\alpha,\beta)\mapsto\frac{e^{2{\rm i}k\alpha}e^{2{\rm i}\ell\beta}}{\pi}\big\}_{(k,\ell)\in\mathbb{Z}^{2}}~.$ (A.7)
@@
$\tilde{p}(\alpha,\lambda)=\sum_{k\in\mathbb{Z}}\,c_{k}(\lambda)\frac{e^{2{\rm i}k\alpha}}{\sqrt{\pi}},~~~\text{where }~~~c_{k}(\lambda)=\int_{0}^{\pi}\tilde{p}(\alpha,\lambda)\frac{e^{-2{\rm i}k\alpha}}{\sqrt{\pi}}{\rm d}\alpha~,$ (A.8)
@@
$\cos\big(2(\alpha\!-\!\beta)\big)=\frac{\pi}{2}\frac{e^{2{\rm i}\alpha}e^{-2{\rm i}\beta}}{\pi}+\frac{\pi}{2}\frac{e^{-2{\rm i}\alpha}e^{2{\rm i}\beta}}{\pi}$ (A.9)
@@
$\begin{array}[]{lll}\int_{\Lambda}\tilde{p}(\alpha,\lambda){\rm d}\rho(\lambda)=\sum_{k\in\mathbb{Z}}\Big(\int_{\Lambda}c_{k}(\lambda){\rm d}\rho(\lambda)\Big)\frac{e^{2{\rm i}k\alpha}}{\sqrt{\pi}}=0,&&(ii)\\[5.69054pt] \sum_{(k,\ell)\in\mathbb{Z}^{2}}\Big(\int_{\Lambda}c_{k}(\lambda)c_{l}(\lambda){\rm d}\rho(\lambda)\Big)\frac{e^{2{\rm i}k\alpha}e^{2{\rm i}\ell\beta}}{\pi}~=~\frac{\pi}{2}\frac{e^{2{\rm i}\alpha}e^{-2{\rm i}\beta}}{\pi}+\frac{\pi}{2}\frac{e^{-2{\rm i}\alpha}e^{2{\rm i}\beta}}{\pi}.&&(iii)\end{array}$ (A.10)
@@
$\begin{array}[]{rll}c_{0}(\lambda)\in\mathbb{R},~~~\forall k\in\mathbb{Z}:&\int_{\Lambda}c_{k}(\lambda){\rm d}\rho(\lambda)&=0.\\[5.69054pt] &\int_{\Lambda}c_{1}(\lambda)c_{-1}(\lambda){\rm d}\rho(\lambda)&=\frac{\pi}{2}\\[5.69054pt] \forall(k,\ell)\in\mathbb{Z}^{2}\backslash\{(1,-1),(-1,1)\}:&\int_{\Lambda}c_{k}(\lambda)c_{l}(\lambda){\rm d}\rho(\lambda)&=0\\[5.69054pt] \forall k\in\mathbb{Z}\backslash\{-1,1\}:&\int_{\Lambda}c_{k}(\lambda)c_{-k}(\lambda){\rm d}\rho(\lambda)&=0~.\end{array}$ (A.11)
@@
$\tilde{p}(\alpha,\lambda)=\dfrac{1}{\sqrt{\pi}}\big(c_{1}(\lambda)e^{2{\rm i}\alpha}+c_{-1}(\lambda)e^{-2{\rm i}\alpha}\big)~=~\dfrac{2|c_{1}(\lambda)|}{\sqrt{\pi}}\cos\big(2\alpha+\alpha(\lambda)\big)~,$ (A.13)
where $\alpha(\lambda)$ is given by $c_{1}=|c_{1}(\lambda)|e^{-i\alpha(\lambda)}$. The values of $c_{\pm 1}(\lambda)$ depend on the choice for $\Lambda$. For example, if $\Lambda=[0,1]\subset\mathbb{R}$ then condition (A.12) is met by $c_{\pm 1}(\lambda)=\sqrt{\frac{\pi}{2}}e^{\pm 2\pi{\rm i}\lambda}$. Note that $\tilde{p}(\alpha,\lambda)$ assumes many more values than just $\pm 1$ as was established in Step 2. With this contradiction in the derived properties of $\tilde{p}(\alpha,\lambda)$ we conclude that $p(x,y,\alpha,\beta)$ is not Bell separable.